OpenStudy (anonymous):

Find the polar form of the equation x^2+y^2=10y

OpenStudy (anonymous):

Okay so the easiest part is: \[ r^2 = x^2 + y^2 \]So we get it like this: \[ r^2=10y \]Next we use: \[ y = r\sin\theta \]To get: \[ r^2 = 10r\sin\theta \]Lastly dividing by \(r\) we get: \[ r = 10\sin\theta \]

OpenStudy (anonymous):

The choices are: A. (r=\[\sqrt{10 \sin \theta}\] B. r= 10 sin theta C. r= 100sin^2 theta D. None of the above

OpenStudy (anonymous):

Just use the following equations: \[ \begin{array}{rcl} r^2 &=& x^2+y^2 \\ x &=& r\sin\theta \\ y &=& r \cos\theta \end{array} \]

OpenStudy (anonymous):

I already pretty much gave you an answer on a platter, I'm not picking out a letter. Try to think about the problem a bit.

OpenStudy (anonymous):

Okay I got it thanks

OpenStudy (anonymous):

oops, I swapped the \(x\) and \(y\) on those equations above. stupid open study thing doesn't let you edit.

OpenStudy (anonymous):

For the record \[ \begin{array}{rcl} r^2 &=& x^2+y^2 \\ y &=& r\sin\theta \\ x &=& r \cos\theta \end{array} \]

OpenStudy (anonymous):

Thanks I will keep all of those equations in mind